OpenStudy (anonymous):

tell whether the lines for each pair of equations are parallel perpendicular or neither 7x - 4y= 4 x - 4y= 3

hero (hero):

`1.` Isolate y in each equation so that they have the form y = mx + b `2.` If the slopes are equal to each other, then both lines are parallel. If the product of the slopes of both equations is -1, then the lines are perpendicular. `1.` Equation 1: 7x - 4y = 4 Add 4y to both sides; Subtract 4 from both sides: 7x - 4 = 4y Divide both sides by 4: \[\frac{7x - 4}{4} = y\] \[\frac{7x}{4} - 1 = y\] Equation 2: x - 4y = 3 add 4y to both sides; subtract 3 from both sides: x - 3 = 4y Divide both sides by 4: \[\frac{x - 3}{4} = y\] \[\frac{x}{4} - \frac{3}{4} = y\] `2.` The slopes of both equations are \(m_1 = \frac{7}{4}\) and \(m_2 = \frac{1}{4}\). Clearly their slopes are not the same. Both slopes are positive so their products will not be negative. Thus the lines are neither parallel nor perpendicular.

OpenStudy (anonymous):

@Hero Thank you so much